# Aioli: A unified optimization framework for language model data mixing

### v1

@@ 4.1 Experiment Details
Data settings.
We use a sampled version of SlimPajama [53, 73], a pre-processed version of the RedPajama pretraining dataset [62], which has been used to train open-source LMs [24, 59]. SlimPajama consists of $7$ data groups: ArXiv, Books, CommonCrawl, C4 [50], Github, StackExchange, and Wikipedia. To develop a fine-grained understanding of data mixing, we create 6 settings by extracting combinations of these groups. We study three settings with $m=2$: Arxiv/StackExchange, Github/C4, and Book/StackExchange. We study two settings with $m=3$: Arxiv/Book/StackExchange and CommonCrawl/Github/Wikipedia. Finally, we study mixing over the full SlimPajama dataset with $m=7$.
Models.
We train 160M parameter GPT-style decoder-only LMs with batch size $8$ and context length $2048$. For $m=2,3$, we train for $5$K steps, and for $m=7$, we train for $40$K steps.
Training sweeps.
To assess the true loss-proportion relationship and compare it to the assumptions made by existing methods, we conduct training sweeps over a set of various mixture proportions, denoted as $\mathcal{P}$. For $m=2$, we set $\mathcal{P}=\{[0.1,0.9],[0.2,0.8],\dots,[0.9,0.1]\}$. For $m=3$ and $7$, we set $\mathcal{P}$ equal to $10$ $\bm{p}$’s drawn from the Dirichlet distribution with $\alpha=1.0$ and $1.5$, respectively.
@@ 4.2 Mixing law parameterization
We examine whether existing methods’ mixing law parameterizations—log-linear static and linear dynamic—capture the true loss-proportion relationship. By empirically fitting them to loss-proportion pairs, we find that both parameterizations are indeed well-specified. Full results for both mixing laws are in Table 5 in Appendix C.1. We also discuss the generality of these parameterizations across training scales and other SlimPajama subsets in Appendix C.1.1, and study if these parameterizations hold over mixtures of instruction-tuning tasks in Appendix C.1.2.
Setup.
For the log-linear static mixing law, we study if there exists $A,b,c$ such that $L_{\text{val},i}(\bm{p})$ can be expressed as $c_{i}+b_{i}\exp(\sum_{j=1}^{m}-A_{ij}p_{j})$ for all $i\in[m]$. We fit the parameters using full training runs on $\mathcal{P}$. For the linear dynamic mixing law, we study if there exists $A^{t}$ such that $L_{\text{val},i}^{t+1}(\bm{p})$ can be expressed as $L_{\text{val},i}^{t}(\bm{p})-\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$, for all $i\in[m]$ ($b^{t}$ is absorbed into $A^{t}$). To fit $A^{t}$, we select a timestep $t$ and train on a static proportion $p^{0}\in\mathcal{P}$ until round $t$, and at $t+1$ we sweep the values of $p^{t+1}\in\mathcal{P}$.
Results.
On average across our $6$ data settings, the mean squared error (MSE) of the fitted log-linear static mixing law is $8.9\times 10^{-4}$, and the $R^{2}$ coefficient of determination is 0.991. The average MSE of the fitted linear dynamic mixing law is $1.0\times 10^{-4}$ and the $R^{2}$ is 0.947. See Figure 2 for examples. Since both parameterizations have high $R^{2}$ and low MSE, we conclude that they capture the true loss-proportion relationship well and are of high fidelity.
@@ 4.3 Values of mixing law parameters
As shown in Table 1, each method sets the parameters of its mixing law differently. We study how close the method-specific parameters are to the optimal parameters that are obtained when fitting the method’s mixing law to the true loss-proportion relationship, and if these parameter disparities are reflected in method performance. We find that existing methods’ differences in parameters are largely responsible for their performance. We omit studying DML since its parameters are fitted from full training runs and hence differ from the optimal parameters in estimation error only.
Setup.
For each online method—Skill-It, DoReMi, and DoGE—we select a step $t$ and obtain the method-specific $A^{t}$. We then sweep $\mathcal{P}$ for the next round $t+1$. This sweep is used to fit an approximately optimal $A^{t\star}$ that captures the true loss-mixture relationship, $L_{\text{val}}^{t+1}(\bm{p})=L_{\text{val}}^{t}(\bm{p})-A^{t\star}p^{t}$, as well as fit a $b^{t}\in\mathbb{R}$ used for scaling $A^{t}$ (details in Appendix C.2). We study the relationship between $\tilde{A}^{t}:=b^{t}A^{t}$ and $A^{t\star}$, and how this relationship is connected to the performance of the method.
@@
To express similarity between $\tilde{A}^{t}$ and $A^{t\star}$ in a way that is reflected in performance, we observe that from Lemma 1, $p^{t}$ is updated using the column sum of $\tilde{A}^{t}$, $\mathbf{1}^{\top}\tilde{A}^{t}$. Moreover, the magnitude of $\tilde{A}^{t}$ is not critical to performance since the step size $\eta$ can always be tuned to control this. Therefore, we compare the vectors $\tilde{a}^{t}=\mathbf{1}^{\top}\tilde{A}^{t}/\|\mathbf{1}^{\top}\tilde{A}^{t}\|_{2}$ and $a^{t\star}=\mathbf{1}^{\top}A^{t\star}/\|\mathbf{1}^{\top}A^{t\star}\|_{2}$. Finally, we note that the order of the elements of $\tilde{a}^{t}$ determines the update direction from $p^{t}$ to $p^{t+1}$ in Lemma 1. Therefore, we propose a similarity score that is an average of cosine similarity and the Spearman rank correlation, $\text{sim}(\tilde{A}^{t},A^{t\star})=0.5\text{cossim}(\tilde{a}^{t},a^{t\star})+0.5\text{Spearman}(\tilde{a}^{t},a^{t\star})$. This metric is bounded between $-1$ and $1$, where $1$ indicates $\tilde{a}^{t}=a^{t\star}$ and $-1$ indicates $\tilde{a}^{t}=-a^{t\star}$.
Results.
In Figure 3, we plot each method’s $\text{sim}(\tilde{A}^{t},A^{t\star})$ versus each method’s improvement over the stratified sampling baseline, which sets $p_{i}=1/m$ for all $i\in[m]$, for each dataset in the $m=2,3$ data settings. We find that no existing online method works well across all datasets (also see Table 2), and that our metric and loss improvement have a moderate positive correlation ($R^{2}=0.491$). This suggests that $A^{t}$’s fidelity to $A^{t\star}$ clearly influences the performance of online methods, and that existing methods’ parameters are not consistently accurate across the datasets. In Appendix C.2.1, we give more details on the properties of $A^{t\star}$, such as how $A^{t\star}$ varies with $t$ and the significance of its off-diagonal entries. We find that restricting $A^{t}$ to be static or diagonal—as some existing methods’ parameters are structured—can result in lower $\text{sim}(\tilde{A}^{t},A^{t\star})$.
@@ 4.4 Solving strategy
We study the assumptions made in how existing methods solve the LMO optimization problem. We find that the greedy approximation used by EGD, $\text{minimize}_{p^{t}}\sum_{i=1}^{m}L_{\text{val},i}^{t+1}(\bm{p})$, does not significantly compromise performance compared to full optimization of dynamic proportions, which has an exponentially large solution space. In particular, we study if greedily selecting $p^{t}$ from $\mathcal{P}$ at each $t$ yields the optimal dynamic proportions in $(\mathcal{P})^{T}$, and we find that this holds in $2$ out of $3$ data settings (Table 10). This suggests that the greedy approximation can simplify optimization without substantial performance loss. We also comment on other possible solving strategies in Appendix C.3.
@@ 5 Aioli: a Method for Improved Data Mixing
Estimating $A^{t\star}$
To build intuition, note that for each group $i\in[m]$, the $m$ variables $A_{i1}^{t},\dots,A_{im}^{t}$ can be learned by solving a system of $m$ equations, each of the form $\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}=L_{\text{val},i}^{t}(\bm{p})-L_{\text{val},i}^{t+1}(\bm{p})$ and with different $\bm{p}$. That is, we could do a training sweep of $m$ different proportions $p^{t,1},p^{t,2},\dots,p^{t,m}\in\triangle^{m}$ and observe how $L_{\text{val},i}^{t+1}$ changes to obtain $A_{i1}^{t},\dots,A_{im}^{t}$. Aioli uses this intuition while avoiding training sweeps. In LearnParams (Algorithm 2) at round $t$, we define $p^{t,i}=(1-\varepsilon)\mathbf{1}_{i}+\varepsilon\text{Unif}(m)$ for all $i\in[m]$ to be a one-hot vector with probability mass $\varepsilon$ distributed uniformly across all groups. We allocate an initial $\delta\in[0,1]$ fraction of each round for estimating $A^{t}$ (i.e., $\delta S/T$ steps), and we determine a randomly interleaved order for training on $p^{t,1},\dots,p^{t,m}$ over this duration. Each time after we train according to some $p^{t,i}$, we record the resulting changes in loss per group. We then average the loss changes per $p^{t,i}$ across the $\delta S/T$ steps, effectively simulating a training sweep at each round without needing $m$ times the compute and having to roll back training to the beginning of the round. Finally, we solve the linear system of equations to obtain $A^{t}$.
Aioli.
In each round of training, we estimate $A^{t}$ using LearnParams and then normalize the entries of $A^{t}$, producing $\bar{A}^{t}$. Otherwise, $A^{t}$ decreases along with loss over time, resulting in the first few $p^{t}$ updates being much larger in magnitude than others. We update the proportions using $p_{j}^{t}\propto p_{j}^{t-1}\exp(\eta\sum_{i=1}^{m}\bar{A}_{ij}^{t})$, and train for the remainder of that round using $p_{j}^{t}$.
Finally, we design Aioli so that it can be used to improve other data mixing methods, which we study in Section 6.2. Mixture proportions can be updated using Aioli either from the start of training or from the middle of a run. In the latter case, we denote an initial static mixture $\bm{p}^{\text{init}}\in\triangle^{m}$ and initial number of steps $S_{\text{init}}$. If $S_{\text{init}}$ is nonzero, Aioli trains according to $\bm{p}_{\text{init}}$ for the first $S_{\text{init}}$ steps before beginning to update the mixture proportions. Aioli is presented in Algorithm 1.
@@
3:Set $\beta=0_{m,m}$.
@@
5:Randomly shuffle $k$ instances of each $i\in[m]$ to create an order $\mathcal{I}\in[m]^{K}$ over indices of data groups.
6:for $k=1,\dots,K$ do
7:     Let $j=\mathcal{I}_{k}$. Train model on mixture $p^{t,j}$ of $D_{\text{train}}$ for one time segment, obtain $f^{t+k\delta/K}$.
@@
9:         Update $\beta_{ij}\leftarrow\beta_{ij}+L_{\text{val},i}(f^{t+(k-1)\delta/K})-L_{\text{val},i}(f^{t+k\delta/K})$ with loss difference on $D_{\text{val}}^{i}$.
@@ 6 Experimental Results
We evaluate all methods in the LMO framework, including Aioli, in two settings. First, we consider an unrestricted additional training budget setting to assess how Aioli compares to other methods in their original form, since each method uses a different number of extra training runs to learn proportions (Section 6.1). Second, we consider a restricted training budget setting to assess if Aioli can enhance existing methods in practical, budget-constrained conditions, where existing methods have less than a full training run to learn mixing proportions (Section 6.2). Hyperparameters and experimental details are available in Appendix D, and ablations and downstream evaluation are in Appendix E.
Data settings and models.
We use the same data settings and models as in Section 4.1, where we train for $S=5$K steps for $m=2,3$-group settings and $S=40$K steps for the full SlimPajama.
Baselines and evaluation.
We consider three online methods (Skill-It, DoGE, DoReMi) and one offline method (DML). We also consider grid search (GS), which sweeps training runs and selects $\bm{p}$ with the lowest average validation loss, and stratified sampling, which sets $p_{i}=\frac{1}{m}$ for all $i\in[m]$. For each method, we report the average test perplexity per group of the trained model. This metric is considered a proxy for downstream performance [21] and also represents the objective in the data mixing problem presented in Section 2.
@@ 6.1 Unrestricted Setting
Setup.
We allow methods up to $10S$ additional training steps to learn the mixture proportions. Approaches like grid search and DML can use the entire budget (searching and fitting over $10$ full runs), while Skill-It, DoReMi, and DoGE use $mS$, $2S$, and $S$ extra training steps, respectively (see Section 3.3). Stratified sampling and Aioli use no extra training steps. We evaluate Aioli with $S_{\text{init}}=0$.
Results.
In Table 2, we find that Aioli robustly outperforms stratified sampling in all 6 data settings by an average of 0.280 perplexity points, while all other methods do worse than stratified sampling on at least 1 set of data groups by up to 6.9 points. The performance of Aioli and other online methods is additionally reflected in Figure 3, in which we find that Aioli’s $A^{t}$ similarity with $A^{t\star}$ is correlated with performance. While Aioli’s parameter similarity is not always the highest, we note that its lowest similarity score is much higher than that of other methods, providing evidence that Aioli’s parameter estimation procedure is more consistently accurate than that of other methods. Lastly, regarding offline methods, we hypothesize that their poor performance on settings with larger $m$ is due to the training budget being limited to $10S$, and that increasing this budget would eventually allow them to perform well.

### v2

@@ 4.1 Experiment Details
Data settings. We use a sampled version of SlimPajama [54, 73], a pre-processed version of the RedPajama pretraining dataset [62]. SlimPajama consists of $7$ data groups: ArXiv, Books, CommonCrawl, C4 [51], Github, StackExchange, and Wikipedia. To develop a fine-grained understanding of data mixing, we create 6 settings by extracting combinations of these groups. We study three settings with $m=2$: Arxiv/Stackexchange, Github/C4, and Book/StackExchange. We study two settings with $m=3$: Arxiv/Book/StackExchange and CommonCrawl/Github/Wikipedia. Finally, we study mixing over the full SlimPajama dataset with $m=7$.
Models. We train 160M parameter GPT-style decoder-only LLMs with batch size $8$ and context length $2048$. For $m=2,3$, we train for $5$K steps, and for $m=7$, we train for $40$K steps.
Training sweeps. To assess the true loss-proportion relationship and compare it to the assumptions made by existing methods, we conduct training sweeps over different mixture proportions, denoted as $\mathcal{P}$. For $m=2$, we set $\mathcal{P}=\{[0.1,0.9],[0.2,0.8],\dots,[0.9,0.1]\}$. For $m=3$ and $7$, we set $\mathcal{P}$ equal to $10$ $\bm{p}$’s and $40$ $\bm{p}$’s drawn from the Dirichlet distribution with $\alpha=1.0$ and $1.5$, respectively.
@@ 4.2 Mixing law parameterization
We examine whether existing methods’ mixing law parameterizations—log-linear static and linear dynamic—capture the true loss-proportion relationship. By empirically fitting them to loss-proportion pairs, we find that both parameterizations are indeed well-specified. Full results for both mixing laws are in Table 5 in Appendix C.1. We discuss the generality of these parameterizations across training scales and other datasets, as well as higher-order parameterizations, in Appendix C.1.1.
Setup. For the log-linear static mixing law, we study if there exists $A,b,c$ such that $L_{\text{val},i}(\bm{p})$ can be expressed as $c_{i}+b_{i}\exp(\sum_{j=1}^{m}-A_{ij}p_{j})$ for all $i\in[m]$. We fit the parameters using full training runs on $\mathcal{P}$. For the linear dynamic mixing law, we study if there exists $A^{t}$ such that $L_{\text{val},i}^{t+1}(\bm{p})$ can be expressed as $L_{\text{val},i}^{t}(\bm{p})-\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t}$, for all $i\in[m]$ ($b^{t}$ is absorbed into $A^{t}$). To fit $A^{t}$, we select a timestep $t$ and train on a static proportion $p^{0}\in\mathcal{P}$ for all $p^{1},\dots,p^{t}$ until time $t$, and at $t+1$ we sweep the values of $p^{t+1}\in\mathcal{P}$.
Results. On average across our $6$ data settings, the mean squared error (MSE) of the fitted log-linear static mixing law is $8.9\times 10^{-4}$, and the $R^{2}$ coefficient of determination is 0.991. The average MSE of the fitted linear dynamic mixing law is $1.0\times 10^{-4}$ and the $R^{2}$ is 0.947. See Figure 2 for examples. Since both parameterizations have high $R^{2}$ and low MSE, we conclude that they capture the true loss-proportion relationship well and are of high fidelity.
@@ 4.3 Values of mixing law parameters
As shown in Table 1, each method sets the parameters of its mixing law differently. We study how close the method-specific parameters are to the optimal parameters that are obtained when fitting the method’s mixing law to the true loss-proportion relationship, and if these parameter disparities are reflected in method performance. We find that existing methods’ differences in mixing law parameters are largely responsible for their performance. We omit studying DML since its parameters are fitted from full training runs and hence differ from the optimal in estimation error only.
@@
Setup. For Skill-It, DoReMi, and DoGE, we select a step $t$ and obtain the method-specific $A^{t}$. We then sweep $\mathcal{P}$ for the next round $t+1$. This sweep is used to approximate an optimal $A^{t\star}$ that captures the true loss-mixture relationship, $L_{\text{val}}^{t+1}(\bm{p})=L_{\text{val}}^{t}(\bm{p})-A^{t\star}p^{t}$, as well as fit a $b^{t}\in\mathbb{R}$ used for scaling $A^{t}$ (details in Appendix C.2). We study the relationship between $\tilde{A}^{t}:=b^{t}A^{t}$ and $A^{t\star}$, and how it is related to the performance of the method.
To express similarity between $\tilde{A}^{t}$ and $A^{t\star}$ in a way that is reflected in performance, we observe that from Lemma 1, $p^{t}$ is updated using the column sum of $A^{t}$, $\mathbf{1}^{\top}A^{t}$. Moreover, the magnitude of $A^{t}$ is not critical to performance since the step size $\eta$ can always be tuned to control this. Therefore, we compare the vectors $\tilde{a}^{t}=\mathbf{1}^{\top}\tilde{A}^{t}/\|\mathbf{1}^{\top}\tilde{A}^{t}\|_{2}$ and $a^{t\star}=\mathbf{1}^{\top}A^{t\star}/\|\mathbf{1}^{\top}A^{t\star}\|_{2}$. Finally, we note that the order of the elements of $\tilde{a}^{t}$ determines the update direction from $p^{t}$ to $p^{t+1}$ in Lemma 1. Therefore, we propose a similarity score that is an average of cosine similarity and the Spearman rank correlation, $\text{sim}(\tilde{A}^{t},A^{t\star})=0.5\text{cossim}(\tilde{a}^{t},a^{t\star})+0.5\text{Spearman}(\tilde{a}^{t},a^{t\star})$. This metric is bounded between $-1$ and $1$, where $1$ indicates $\tilde{a}^{t}=a^{t\star}$ and $-1$ indicates $\tilde{a}^{t}=-a^{t\star}$.
Results. In Figure 3, we plot each method’s $\text{sim}(\tilde{A}^{t},A^{t\star})$ versus each method’s improvement over the stratified sampling baseline, which sets $p_{i}=1/m$ for all $i\in[m]$, for each dataset in the $m=2,3$ data settings. We find that no existing online method works well across all datasets (also see Table 2), and that our metric and loss improvement have a moderate positive correlation ($R^{2}=0.491$). This suggests that $A^{t}$’s accuracy is critical to the performance of online methods, and that existing methods’ $A^{t}$ are not consistently accurate across the datasets. In Appendix C.2.1, we give more details on the structure of $A^{t\star}$, providing intuition for why existing methods’ parameters cannot express it.
@@ 4.4 Solving strategy
We study the assumptions made in how existing methods solve the LMO optimization problem. We find that the greedy approximation used by EGD, $\text{minimize}_{p^{t}}\sum_{i=1}^{m}L_{\text{val},i}^{t+1}(\bm{p})$, does not significantly compromise performance compared to full optimization of dynamic proportions, which has an exponentially large solution space. In particular, we study if greedily selecting $p^{t}$ from $\mathcal{P}$ at each $t$ yields the optimal dynamic proportions in $\mathcal{P}^{T}$, and we find that this holds in $2$ out of $3$ data settings (Table 10). This suggests that the greedy approximation can simplify optimization without substantial performance loss. We also comment on other possible solving strategies in Appendix C.3.
@@ 5 Aioli: a Method for Improved Data Mixing
Estimating $A^{t\star}$. To build intuition, we first consider a high-cost naive approach. For each round, we could conduct a training sweep of $m$ different proportions $p^{t,1},\dots,p^{t,m}$, and observe each resulting change in loss. We could then solve a system of $m$ equations for each $i$: $L_{\text{val},i}^{t}-L_{\text{val},i}^{t+1}(p^{t,s})=\sum_{j=1}^{m}A_{ij}^{t}p_{j}^{t,s}$ for $s\in[m]$, obtaining vectors $A_{1}^{t},\dots,A_{m}^{t}$. However, this approach effectively requires $m$ extra training runs.
Aioli similarly solves a system of equations, but it computes loss changes per sweep mixture without requiring extra training. First, it allocates $\delta$ fraction of the training round for learning $A^{t}$. Second, it partitions this $\delta$ into $K=mk$ intervals and trains according to an interleaved order on $p^{t,1},\dots,p^{t,m}$. After training on each $p^{t,j}$, we record the resulting change in validation losses, and we average over all of $p^{t,j}$’s intervals. Intuitively, the interleaving ensures that the model is trained on each $p^{t,j}$ for several intervals throughout $\delta$, which can approximate if we were to train on $p^{t,j}$ for the entire $\delta$ (which approximates the entire round). This procedure is outlined in LearnParams (Alg. 2), with more details in Appendix D and Figure 7.
Aioli. First, we set $p^{0}$ to be uniform. In each round, we estimate $A^{t}$ using LearnParams and then normalize the entries of $A^{t}$, producing $\bar{A}^{t}$. Otherwise, $A^{t}$ decreases along with loss over time, resulting in the first few $p^{t}$ updates being much larger in magnitude than others. Then, we update the proportions using $p_{j}^{t}\propto p_{j}^{t-1}\exp(\eta\sum_{i=1}^{m}\bar{A}_{ij}^{t})$, as in Lemma 1, and train for the remainder of that round using $p_{j}^{t}$.
Finally, we design Aioli so that it can also be used to improve other data mixing methods, which we study in Section 6.2. Mixture proportions can be updated using Aioli either from the start of training or from the middle of a run. In the latter case, we denote an initial static mixture $\bm{p}^{\text{init}}\in\triangle^{m}$ and initial number of steps $S_{\text{init}}$. If $S_{\text{init}}$ is nonzero, Aioli trains according to $\bm{p}_{\text{init}}$ for the first $S_{\text{init}}$ steps before updating the mixture proportions. Aioli is presented in Algorithm 1.
@@
2:Split the fraction of a training round $\delta$ into $K$ intervals, where $K=mk$.
3:Set $\beta=0_{m,m}$
@@
5:Randomly shuffle $k$ instances of each $i\in[m]$ to create an order $\mathcal{I}\in[m]^{K}$.
6:for $\tau=1,\dots,K$ do
7:     Let $j=\mathcal{I}_{\tau}$. Train model on mixture $p^{t,j}$ of $D_{\text{train}}$ for one interval, obtain $f^{t+\tau\delta/K}$.
@@
9:         Update $\beta_{ij}\leftarrow\beta_{ij}+L_{\text{val},i}(f^{t+(\tau-1)\delta/K})-L_{\text{val},i}(f^{t+\tau\delta/K})$ with loss difference on $D_{\text{val}}^{i}$.
@@ 6 Experimental Results
We evaluate all methods in the LMO framework, including Aioli, in two settings. First, we consider an unrestricted additional training budget setting to assess how Aioli compares to other methods in their original form, since each method uses a different number of extra training runs to learn proportions (Section 6.1). Second, we consider a restricted training budget setting to assess if Aioli can enhance existing methods in practical, budget-constrained conditions, where existing methods have less than a full training run to learn mixing proportions (Section 6.2). Hyperparameters and experimental details, including proportion trajectories are available in Appendix E. Downstream evaluation, ablations, experiments on larger models, and results adapting Aioli to an out-of-domain setting are in Appendix F.
Data settings and models. We use the same data settings and models as in Section 4.1, where we train for $S=5$K steps for $m=2,3$-group settings and $S=40$K steps for the full SlimPajama.
Baselines and evaluation. We consider three online methods (Skill-It, DoGE, DoReMi) and one offline method (DML). We also consider grid search (GS), which sweeps training runs and selects $\bm{p}$ with the lowest average validation loss, and stratified sampling, which sets $p_{i}=\frac{1}{m}$ for all $i\in[m]$. For each method, we report the average test perplexity per group of the trained model. This metric is considered a proxy for downstream performance [22] and also represents the objective in the data mixing problem.
@@ 6.1 Unrestricted Setting
Setup. We allow methods up to $10S$ additional training steps to learn the mixture proportions. Approaches like grid search and DML can use the entire budget (searching and fitting over $10$ full runs), while Skill-It, DoReMi, and DoGE use $mS$, $2S$, and $S$ extra training steps, respectively (see Section 3.3). Stratified sampling and Aioli use no extra training steps. We evaluate Aioli with $S_{\text{init}}=0$.
Results. In Table 2, we find that Aioli robustly outperforms stratified sampling in all 6 data settings by an average of 0.274 perplexity points, while all other methods do worse than stratified sampling on at least 1 set of data groups by up to 6.9 points. The performance of Aioli and other online methods is additionally reflected in Figure 3, in which we find that Aioli’s $A^{t}$ similarity with $A^{t\star}$ is correlated with performance. While Aioli’s parameter similarity is not always the highest, we note that its lowest similarity score is much higher than that of other methods, providing evidence that Aioli’s parameter estimation procedure is more consistently accurate than that of other methods. Lastly, regarding offline methods, we hypothesize that their poor performance on settings with larger $m$ is due to the training budget being limited to $10S$, and that increasing this budget would eventually allow them to perform well.